OpenStudy (anonymous):

how do you solve a quadratic equation using the quadratic formula. Can someone explain this and show me an example so i get it...

OpenStudy (anonymous):

\[ax^2 + bx +c = 0 \implies x=\frac{-(b) \pm \sqrt{b^2 - 4(a)(c)}}{2(a)}\]

myininaya (myininaya):

did you look at that scan i made for you?

OpenStudy (anonymous):

Just plug them in

OpenStudy (anonymous):

yes...I could not hardly read it.. sorry

myininaya (myininaya):

ok nvm then

OpenStudy (anonymous):

2x^2 + 10x + 3 = 0 a = 2 b = 10 c =3 D = sqrt(b^2-4ac) = sqrt (100-24) = sqrt(76) x = -b +- sqrt (D) ------------------ 2a

OpenStudy (anonymous):

So basically if you get something of the form you have on the right you plug it in to the formula on the left and you find the x values that make the right hand equation true.

OpenStudy (anonymous):

Err.. I have a hard time with right/left. So reverse what I said ;)

OpenStudy (anonymous):

Thank you..lol

myininaya (myininaya):

i hate right and left i use postive and negative when i talk

myininaya (myininaya):

i say lets go in the positive direction hoping people think i mean is it right?

OpenStudy (anonymous):

I appreciate all the help

myininaya (myininaya):

or left?

OpenStudy (anonymous):

ok

myininaya (myininaya):

i think its right

myininaya (myininaya):

polpak can i challenge you?

myininaya (myininaya):

lol

OpenStudy (anonymous):

with?